# A Scalable Pipeline for Enabling Non-Verbal Speech Generation and Understanding

### v1

@@ Abstract
Human spoken communication involves not only lexical content but also non-verbal vocalizations (NVs) such as laughter, sighs, and coughs, which convey emotions, intentions, and social signals. However, most existing speech systems focus solely on verbal content and lack the ability to understand and generate such non-verbal cues, reducing the emotional intelligence and communicative richness of spoken interfaces. In this work, we introduce NonVerbalSpeech-38K, a large and diverse dataset for non-verbal speech generation and understanding, collected from real-world media and annotated using an automatic pipeline. The dataset contains 38,718 samples (about 131 hours) with 10 categories of non-verbal cues, such as laughter, sniff, and throat clearing. We further validate the dataset by fine-tuning state-of-the-art models, including F5-TTS and Qwen2-Audio, demonstrating its effectiveness in non-verbal speech generation and understanding tasks. Our contributions are threefold: (1) We propose a practical pipeline for building natural and diverse non-verbal speech datasets; (2) We release a large-scale dataset to advance research on non-verbal speech generation and understanding; (3) We validate the dataset’s effectiveness by demonstrating improvements in both non-verbal speech synthesis and captioning, thereby facilitating richer human-computer interaction.
@@ 1 Introduction
Spoken interaction has become a core interface in human-computer communication, driven by the rapid advancement of automatic speech recognition (ASR), text-to-speech (TTS) synthesis, and large language models (LLMs). These technologies have enabled systems to understand and generate semantically coherent spoken content with impressive fluency (Zhang et al. 2023; Xie and Wu 2024; Fang et al. ; Fang et al. 2024; Défossez et al. 2024). However, human communication is far richer than just words. Non-verbal vocalizations (NVs), such as laughter, sighs and other acoustic events not tied to lexical meaning, are crucial for conveying emotion, intentions, and social signals (Truong and Van Leeuwen 2007; Cortes et al. 2021; Cowen et al. 2019; Kamiloğlu, Fischer, and Sauter 2020; Lima et al. 2014). Despite their importance, most current speech systems lack the ability to perceive or express such non-verbal elements, a limitation that hinders the creation of truly human-like and empathetic interactions.
@@
For an ideal speech system to genuinely mimic human communication, it must be capable of both understanding and generating non-verbal sounds. In this work, we focus on two key tasks that reflect these capabilities: non-verbal speech generation and non-verbal speech understanding (Figure 1).
@@
Speech synthesis with NVs can be categorized into two distinct approaches. The first is non-verbal tag-controlled TTS, which inserts explicit tags (e.g., [laughing], [breath]) into text to control the type and position of NVs (Guo et al. 2024; Du et al. 2024a, b, 2025; 2noise 2024; Nari-labs 2025). These methods offer fine-grained control but require annotated training data and still struggle with precise timing and occurrence control. The second is spontaneous-style TTS, which predicts NVs directly from contextual cues without requiring aligned annotations (Li et al. 2024a; 2noise 2024). While more flexible and easier to scale, such models often rely on proprietary data or suffer from limited data quality, hindering reproducibility and generalization.
Figure 1: Task definition of non-verbal TTS and non-verbal speech understanding.
@@
Beyond speech synthesis, understanding non-verbal vocalizations is equally important for building truly human-like speech interaction systems. Traditionally, speech processing systems have focused on single-domain tasks such as automatic speech recognition (ASR) or sound event classification independently. However, in real-world scenarios, speech and non-verbal sounds naturally co-occur, driving a shift toward unified models that can handle both simultaneously. Recent speech understanding models  (Tang et al. 2024; Chu et al. 2023, 2024; Huang et al. 2025; Wu et al. 2025; Ding et al. 2025) employ multi-task training strategies to jointly perform ASR, sound event classification, and related speech tasks. Although this unified approach integrates multiple tasks within a single framework, effectively combining these capabilities—particularly transcribing non-verbal information embedded within speech—remains challenging (Shione, Wakabayashi, and Kitaoka 2023b, a). Progress is further limited by the scarcity of large-scale, fine-grained annotated datasets.
As outlined above, the development of both non-verbal speech synthesis and understanding remains fundamentally hindered by the scarcity of publicly available, large-scale, diverse, and well-annotated datasets. To address this, several efforts have attempted to incorporate non-verbal vocalizations (NVs) into speech datasets (Xin, Takamichi, and Saruwatari 2024; Xin et al. 2024; Wang et al. 2025; Nguyen et al. 2023). However, existing datasets often face significant limitations. Some works rely on manual annotation or human recording, which severely restricts speaker diversity and the scale of the dataset. Other approaches use rule-based data augmentation, which may lead to issues with speech naturalness and timbre inconsistency. Additionally, recent NonverbalTTS (Borisov, Spirin, and Diatlova 2025) has explored semi-supervised pipeline to increase scale, but the resulting dataset is still limited in size and coverage.
Figure 3: Data preparation and frame-level detection model training.
To address these limitations, we present NonVerbalSpeech-38K, a large-scale dataset of naturally occurring non-verbal vocal expressions with rich text annotations. The data is collected from diverse real-world sources, including animations, movies, variety shows, and radio dramas (Table 3). To ensure scalability, we develop an automatic annotation pipeline for detecting non-verbal events and integrating them with speech transcriptions (Figure 2). The resulting dataset contains 38,718 samples (approximately 131 hours), covering 10 types of non-verbal tags such as cough and laugh (Figure 4). Table 1 highlights the key differences between our dataset and existing resources. We validate the effectiveness of our dataset by fine-tuning two representative models: F5-TTS (Chen et al. 2024) for non-verbal TTS, and Qwen-Audio2 (Chu et al. 2024) for end-to-end captioning that incorporates both verbal and non-verbal cues.
Our main contributions are summarized as follows:
  • We design a scalable pipeline for accurate detection and labeling of non-verbal expressions from diverse real-world audio.
  • We release a large-scale annotated dataset to facilitate research on non-verbal TTS and non-verbal speech understanding.
  • We fine-tune F5-TTS to demonstrate the dataset’s utility for non-verbal speech synthesis.
  • We fine-tune Qwen2-Audio to investigate joint modeling of speech transcription and non-verbal tag prediction within an end-to-end captioning framework.
2 Related Work
2.1 Speech Dataset with NVs
Recent speech datasets have begun to include non-verbal annotations. For example, JNV corpus (Xin, Takamichi, and Saruwatari 2024) constructs a Japanese NVs dataset from 4 speakers, covering 6 emotions. Then, JVNV (Xin et al. 2024) further uses these NV phrases to create emotional scripts with ChatGPT (Hurst et al. 2024). Besides, Expresso (Nguyen et al. 2023) captures non-verbal vocalizations from natural expressive speech. It incorporates NVs as one of 26 expressive styles, covering a broad range of vocal behaviors that are inherently difficult to transcribe. Despite their valuable resources for non-verbal speech, their limited speakers count severely restricts data diversity. To broaden coverage of NVs and improve diversity, Capspeech-SE (Wang et al. 2025) constructs a non-verbal speech dataset via a rule-based method that integrates multi-speaker speech with sound effect clips. While this creates a large-scale and diverse dataset, it suffers from limitations in timbre consistency and speech naturalness, which hinder its suitability for TTS tasks. Extending previous efforts, NonverbalTTS (Borisov, Spirin, and Diatlova 2025) creates a 17-hour dataset by augmenting VoxCeleb (Nagrani, Chung, and Zisserman 2017; Chung, Nagrani, and Zisserman 2018) and Expresso (Nguyen et al. 2023) with diverse NVs and emotional annotations through a semi-automated pipeline. Although derived from natural speech, it remains limited in scale, and its pipeline lacks scalability. A comparison of the recently proposed datasets Capspeech (Wang et al. 2025) and NonverbalTTS (Borisov, Spirin, and Diatlova 2025) with our dataset is summarized in Table 1.
Dataset Language #Samples Source Construction Pipeline # Non-Verbal Tags
NonVerbalTTS EN $\sim$6K VoxCeleb and Expresso Semi-automatic 10
Capspeech EN $\sim$37K LibriTTS-R Data augmentation 3
NonVerbalSpeech-38K (Ours) EN, ZH $\sim$38K In-the-wild Automatic 10
Table 1: Comparison of non-verbal speech datasets in terms of language coverage, size, source, construction pipeline, and supported non-verbal tag types. For Capspeech, we include only the samples containing any of the six target non-verbal tags, as shown in Table 5.
Category Count Category Count
Laughing 11,441 Coughing 9,105
Breath 4,851 Sniff 4,564
Crying 4,369 Throat clearing 4,235
Sigh 2,069 Snore 731
Gasp 333 Yawn 266
Table 2: Statistics of non-verbal segments for training the frame-level detection model.
2.2 Speech Synthesis with NVs
Research on speech synthesis with NVs falls into two main approaches. The first is non-verbal tag-controlled TTS, which inserts explicit tags into the input text to specify NV type and position, such as [laughing] or [breath]. By training with NV-labeled speech data, FireRedTTS (Guo et al. 2024) and CosyVoice (Du et al. 2024b) have successfully generated natural speech with non-verbal vocalizations. Dia (Nari-labs 2025) further extends this capability to double-speaker dialogue scenarios. The second category is spontaneous-style TTS, which predicts non-verbal vocalizations directly from contextual cues, thus eliminating the need for explicit annotations. Examples include SponTTS (Li et al. 2024a) and ChatTTS (2noise 2024). However, these models often depend on proprietary datasets or are limited by the scale and quality of publicly available data, restricting their generalization and further development.
2.3 Speech Understanding with NVs
Traditional speech processing systems typically focused on single-domain tasks. Nowadays, the current trend has shifted towards building unified speech understanding models that employ multi-task training strategies to jointly address multi-dimension recognition (Tang et al. 2024; Chu et al. 2023, 2024; Huang et al. 2025; Wu et al. 2025). While this unified approach integrates multiple tasks within a single framework, effectively combining these capabilities—especially the transcription of non-verbal information embedded within spoken content—remains challenging (Shione, Wakabayashi, and Kitaoka 2023b, a).
@@ 3 Methodology
This section presents our data processing pipeline, NonVerbalSpeech-38K Pipeline, as illustrated in Figure 2. It consists of three main stages: Data Preparation and Preprocessing, Non-Verbal Segment Detection, and Integration of Non-Verbal Tags into Speech Content. Among these, Non-Verbal Segment Detection serves as the core component of the pipeline. We detail each step below.
@@ 3.1 Data Preparation and Preprocessing
As shown in Figure 2 (Step 1), we collect about 33,252 hours of data from diverse sources such as radio dramas, cartoons, and comedy sketches (Table 3). Using the Emilia-Pipeline (He et al. 2024), we preprocess the data with standardization, source separation, and speaker diarization to extract single-speaker segments. Voice activity detection (VAD) removes silence, followed by transcription with Whisper-large-v3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022). Segments with WER $20\%$ or DNSMOS <1.0<1.0 are filtered out to ensure high-quality speech. This process also produces segment and word-level timestamps for further processing.
@@ 3.2 Non-Verbal Segments Detection
We train a frame-level detection model to predict precise timestamps of non-verbal events, crucial for embedding non-verbal tags in transcriptions. We first describe the training procedure, then explain its application in the pipeline.
@@
Inspired by T-UAED (Jiang et al. 2025), we propose a frame-level detection model to localize non-verbal events, as illustrated in Figure 3. Specifically, we employ a pretrained wav2vec 2.0 model (Baevski et al. 2020) to extract $L$-layer features $F_{i}\in\mathbb{R}^{B\times T\times D}$. To effectively utilize multi-level representations, we apply a learnable weighted sum:
@@
Figure 4: Language and non-verbal label distribution in the proposed NonVerbalSpeech-38K dataset.
@@
To train the frame-level detection model, we construct augmented samples by concatenating or overlaying non-verbal and speech segments, as shown in Figure 3. Non-verbal clips are collected from widely used sound event datasets including VocalSound (Gong, Yu, and Glass 2022), VGGSound (Chen et al. 2020), Nonspeech7k (Rashid, Li, and Du 2023), AudioSet-Strong, and ESC-50 (Piczak 2015), forming our Non-Verbal Event Database. To ensure quality, we apply energy-based VAD to remove silence. Segments with low CLAP scores (below 0.4) or very short durations (less than 0.3 seconds) are discarded. The resulting high-quality clips are referred to as HQ Non-Verbal Units, with statistics summarized in Table 2. Speech data is drawn from the GigaSpeech Medium subset (Chen et al. 2021) and SEED-TTS-Eval (Anastassiou et al. 2024). Only the GigaSpeech test split and the SEED-TTS-Eval set are used for evaluation.
@@
We set $N_{q}=10$, corresponding to the 10 non-verbal types in Table 2. Both encoder and decoder have 6 layers. The model is fine-tuned using LoRA (Hu et al. 2022) for 300 epochs with a batch size of 64 per GPU on 8 H100 GPUs, using a learning rate of $1\times 10^{-4}$ with warm-up and cosine annealing. After training, the model achieves a frame-level F1-score of around 91% on three test splits. Additional details are provided in the Appendix.
After training the frame-level detection model, we apply it to standardized waveforms using a sliding window strategy, as shown in Figure  2 (Step 2). Standardized waveforms are used because we observe that tools such as Silero-VAD (Team 2024) (used in the Emilia pipeline (He et al. 2024)) often fail to detect non-verbal events occurring at the beginning or end of speech segments. As a result, non-verbal cues at segment boundaries are frequently missed. To provide sufficient context for each frame, we use a window size of 10 seconds and a hop size of 5 seconds.
To improve segment quality, we introduce three filters: first, discard segments shorter than 0.3 seconds (based on empirical observations);); second, compute the maximum energy of each segment and remove those with very low energy, such as subtle breathing sounds that are hard to perceive. In our pipeline, the energy threshold is set to $35\,\mathrm{dB}$. Third, remove segments with a CLAP score (Elizalde et al. 2023) less than 0.3, as a higher CLAP score indicates that the sound truly corresponds to non-verbal events.
3.3 Incorporating Non-Verbal Tags into Speech Content
After filtering out low-quality candidates, we match each non-verbal segment with the nearest speech region obtained from preprocessing. It should be noted that non-verbal events generally do not lie entirely within speech regions. Therefore, any non-verbal segment located more than 1 second away from the nearest speech region is discarded. The pseudocode for identifying the nearest speech region is provided in the Appendix.
After this, we obtain high-quality speech segments with non-verbal events occurring in the current segment. The available information includes specific non-verbal labels with their start and end timestamps, as well as speech transcriptions with word-level timestamps. To support non-verbal speech generation and understanding, it is necessary to integrate the textual content with non-verbal tags. As shown in Step 3 of Figure 2, for a speech segment with the transcription “His funny face made us laugh”, there exists a laughter sound right after the utterance “His funny face”. By aligning the timestamps of the laughter and the speech, we aim to create a final combined transcript like: “His funny face [laugh] made us laugh.”. The related algorithm is provided in the Appendix.
Source
Radio
Dramas
Comedy
Sketches
Cartoon
Variety
Shows
Short
Plays
Crawled (hrs) 17,400 7,200 3,600 1,400 1,200
NVS-38K (#) 21,668 7,273 5,019 1,217 809
Source
Speeches
Docum-
entaries
Movies
Audio-
books
Toy
Unboxing
Crawled (hrs) 1,079 600 500 263 9
NVS-38K (#) 158 105 1,090 1,375 4
Table 3: “Crawled (hrs)” shows the duration of the original crawled data; “NVS-38K (#)” shows the final sample counts in the proposed NonVerbalSpeech-38K dataset.
4 The NonVerbalSpeech-38K Dataset
Using the proposed NonVerbalSpeech-38K pipeline, we construct a dataset from diverse online speech sources, covering a wide range of natural speaking styles and rich non-verbal vocalizations. The raw data statistics are shown in Table 3. Radio dramas, comedy sketches, and cartoons contribute the most data, while movies and audiobooks yield the highest processing conversion rates, suggesting richness in non-verbal content. The initial release contains $\sim$131 hours of English and Chinese speech, totaling about 38,000 samples. As shown in Figure 4, Chinese samples dominate due to the composition of the source data. Interestingly, the frame-level detection model trained only on English still generalizes well to Chinese, indicating strong cross-lingual scalability. Figure 4 also shows the distribution of non-verbal tags. Frequent types such as sigh, sniff, laughing, and coughing dominate, reflecting their prevalence in daily speech. Less frequent types (e.g., crying, yawn) are also included. The resulting durations mostly fall within the range of 3 to 30 seconds (Figure 5). More information is available in the Appendix.
5 Experiments
We evaluate the effectiveness of our dataset through two downstream tasks: non-verbal speech generation and end-to-end non-verbal speech understanding. The details are as follows.

### v2

@@ Abstract
Non-verbal Vocalizations (NVs), such as laughter and sighs, are vital for conveying emotion and intention in human speech, yet most existing speech systems neglect them, which severely compromises communicative richness and emotional intelligence. Existing methods for NVs acquisition are either costly and unscalable (relying on manual annotation/recording) or unnatural (relying on rule-based synthesis). To address these limitations, we propose a highly scalable automatic annotation framework to label non-verbal phenomena from natural speech, which is low-cost, easily extendable, and inherently diverse and natural. This framework leverages a unified detection model to accurately identify NVs in natural speech and integrates them with transcripts via temporal-semantic alignment method. Using this framework, we created and released NonVerbalSpeech-38K, a diverse, real-world dataset featuring 38,718 samples across 10 NV categories collected from in-the-wild media. Experimental results demonstrate that our dataset provides superior controllability for NVs generation and achieves comparable performance for NVs understanding.
Demoshttps://nonverbalspeech38k.github.io/nonverspeech38k/
Datasetshttps://huggingface.co/datasets/nonverbalspeech/nonverbalspeech38k
@@ 1 Introduction
Spoken interaction has become a core interface in human-computer communication, driven by the rapid advancement of automatic speech recognition (ASR), text-to-speech (TTS) synthesis, and large language models. These technologies have enabled systems to generate spoken content that is semantically coherent and impressively fluent, while also accurately understanding semantic input  (Zhang et al. 2023; Xie and Wu 2024; Défossez et al. 2024). However, human communication is far richer than just words. Non-verbal Vocalizations (NVs), such as laughter and sighs, are critical components of human spoken communication, carrying essential emotional, intentional, and social cues beyond lexical content (Truong and Van Leeuwen 2007; Cortes et al. 2021; Cowen et al. 2019; Lima et al. 2014). Despite their importance, most current speech systems lack the ability to perceive or express such NVs, hindering the creation of truly human-like interactions.
@@
To genuinely emulate human communication, a speech system should both recognize and produce NVs. This work focuses on two tasks capturing these capabilities: NV speech generation and understanding (Fig. 1 (a)).
@@
Speech synthesis with NVs can be categorized into two distinct approaches. The first is tag-controlled, which inserts explicit tags (e.g., [laughing], [breath]) into text to control the type and position of NVs (Du et al. 2024; Nari-labs 2025). These methods offer fine-grained control but require annotated training data and still struggle with precise timing and occurrence control. The second is spontaneous-style, which predicts NVs directly from contextual cues (Li et al. 2024). While generally producing more natural results, such models also rely heavily on proprietary annotated data, limiting further research.
Figure 1: (a) The task definition, (b) The proposed frame-level detection model, and (c) The TSA method used in the pipeline (Fig. 2).
Table 1: Comparison of Existing Corpora and NVS.
Dataset Lang. # Utter. Source Method # Types Human?
NonVerbalTTS(Borisov et al. 2025) EN $\sim$6K VoxCeleb, Expresso Manual annotation 10 Y
CapSpeech(Wang et al. 2025) EN $\sim$37K LibriTTS-R Rule-based concatenation / overlay 3 N
SMIIP-NV(Wu et al. 2025b) ZH $\sim$17K - Manual recording 3 Y
MNV-17(Mai et al. 2025) ZH $\sim$3K - Manual recording 17 Y
SynParaSpeech(Bai et al. 2025) ZH $\sim$56K - Rule-based synthesis with manual verification 6 Y
NVSpeech(Liao et al. 2025) ZH $\sim$174K Nonspeech7k,Emilia, Genshin,StarRail Manual labeling followed by model annotation 18 Y
NVS (Ours) EN, ZH $\sim$38K In-the-wild Fully automatic annotation 10 N
  • Unlike existing works, the NVS dataset is built using a fully automated annotation pipeline, requiring no human effort and operating directly on in-the-wild data, which allows for straightforward scaling. “# Types” denotes the number of distinct non-verbal types. “Human?” indicates whether human effort is required.
Beyond speech synthesis, understanding NVs is equally important for building truly human-like speech interaction systems. Traditionally, speech processing systems have focused on single-domain tasks such as ASR or sound event classification independently. However, in real-world scenarios, speech and NV sounds naturally co-occur, driving a shift toward unified models that can handle both simultaneously. Recent speech understanding models  (Tang et al. 2024; Chu et al. 2023, 2024; Wu et al. 2025a; Ding et al. 2025) employ multi-task training strategies to jointly perform ASR, sound event classification, and related speech tasks. Although this unified approach integrates multiple tasks within a single framework, effectively combining these capabilities—particularly transcribing NVs information embedded within speech transcription—remains challenging (Shione et al. 2023b, a). Progress is further constrained by the scarcity of large, well-annotated datasets.
As outlined above, the development of both NV speech synthesis and understanding remains fundamentally hindered by the scarcity of publicly available datasets. To address this, several efforts have attempted to incorporate NVs into speech datasets. However, existing datasets often face significant limitations. Specifically, some works rely on manual annotation or human recording, which severely restricts speaker diversity and the scale of the dataset (Mai et al. 2025; Wu et al. 2025b; Borisov et al. 2025). Other approaches use rule-based data augmentation, which may lead to issues with speech naturalness and timbre inconsistency (Wang et al. 2025; Bai et al. 2025). Recent approaches that first annotate a seed dataset and then train a model for automatic expansion (Liao et al. 2025) still rely heavily on costly human effort for the initial labeling.
To address these limitations, we propose a fully automatic annotation framework (Fig. 2) that annotates NVs from in-the-wild audio without introducing costly human effort. Building on this framework, we introduce NonVerbalSpeech-38K (NVS), a large-scale dataset containing naturally occurring NVs along with NV-aware transcriptions. Collected from diverse in-the-wild sources such as animations, movies, variety shows, and radio dramas, NVS contain 38K samples ($\sim$131 hours) covering 10 NV types, including [coughing] and [laughing], as detailed in Fig. 4. Table 1 highlights the key differences between our dataset and existing resources. The key distinctions lie in three aspects: (1) our dataset supports both English and Chinese, covering a broader linguistic scope; (2) its construction is fully automatic, eliminating costly human annotation; and (3) it is derived directly from in-the-wild audio, allowing the dataset to scale continuously.
We evaluate our dataset by fine-tuning F5-TTS (Chen et al. 2024) for NV synthesis and Qwen2-Audio (Chu et al. 2024) and Whisper-Large-V3 (Radford et al. 2023) for end-to-end transcription of both verbal content and NVs. Results show that our dataset offers better controllability for NV generation and achieves competitive performance in NV speech transcription. Demos are available online [^1] https://nonverbalspeech38k.github.io/nonverspeech38k/.
Our main contributions are as follows:
  • A practical, scalable pipeline with two novel components—a unified NV detection model and a precise TSA module—for automatically and accurately detecting and labeling NV expressions from in-the-wild audio.
  • A large-scale, diverse, well-annotated, real-world dataset covering 10 NV categories for expressive NV TTS and precise NV understanding has been released [^2] https://huggingface.co/datasets/nonverbalspeech/nonverbalspeech38k.
  • Fine-tuning F5-TTS with our NVS (TSA) consistently outperforms the strongest baseline (CLAP score $0.110$ (EN) and $0.100$ (ZH); IMOS $3.22$ (EN) and $0.33$ (ZH)) (Table 2), achieving relative gains of $70\%$, $79\%$, $8\%$, and $11\%$, respectively, showing superior NV controllability.
  • Fine-tuning Qwen2-Audio and Whisper-Large-V3 with our NVS (TSA) reduces TPD (sec. 4.2) by 0.92 and 2.64, respectively, compared with the baseline (Table 4), indicating enhanced NV perception.
@@ 2 Methodology
This section presents our NonVerbalSpeech-38K annotation pipeline (Fig. 2). It consists of three main stages: Data Preparation and Preprocessing, NV Segment Detection, and Incorporation of NV Tags into Speech Content. We describe each stage in detail below.
@@ 2.1 Data Preparation and Preprocessing
As shown in Fig. 2 (Step 1), we first gather in-the-wild data from various sources, such as movies and cartoons. Using the Emilia-Pipeline (He et al. 2024), we preprocess the data through standardization, source separation, and speaker diarization to extract single-speaker segments. Voice activity detection (VAD) removes silence, followed by transcription with Whisper-Large-V3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022). Segments with WER $20\%$ or DNSMOS <1.0<1.0 are filtered to ensure high-quality speech. This process also yields segment- and word-level timestamps for further processing.
@@ 2.2 Non-Verbal Segments Detection
In this step, we first train the unified frame-level NV detection model and then apply it within the pipeline.
@@
Inspired by T-UAED (Jiang et al. 2025), we propose a unified frame-level detection model to localize NV events, as illustrated in Fig. 1 (b). Specifically, we employ a pretrained wav2vec-bert 2.0 model  (Barrault et al. 2023) to extract $L$-layer features $F_{i}\in\mathbb{R}^{B\times T\times D}$. To effectively utilize multi-level representations, we apply a learnable weighted sum:
@@
To train the frame-level detection model, we construct augmented samples by concatenating or overlaying NVs and speech segments (Fig. 1 (b)). NV clips are collected from widely used sound event datasets including VocalSound  (Gong et al. 2022) , VGGSound  (Chen et al. 2020) , Nonspeech7k  (Rashid et al. 2023) , AudioSet-Strong  (Hershey et al. 2021) , and ESC-50  (Piczak 2015) . To ensure quality, we first apply energy-based VAD to remove silence and split each NV clip into multiple segments. Segments with low CLAP scores (<< 0.4) or very short durations (<< 0.3s) are discarded. The resulting high-quality segments are used to augment training samples, with their distribution shown in Fig. 3. Speech data is drawn from the GigaSpeech M subset ($\sim$1,000 hrs) (Chen et al. 2021) and SEED-TTS-Eval (Anastassiou et al. 2024) . The GigaSpeech test split and the SEED-TTS-Eval set are used exclusively for evaluation.
@@
We set $N_{q}=10$, corresponding to the 10 NV types in Fig. 3. Both encoder and decoder have 6 layers. The model is fine-tuned with LoRA (Hu et al. 2022) for 300 epochs on 8 H100 GPUs, using a batch size of 64 per GPU. The learning rate is set to $1\times 10^{-4}$, with a 5% warm-up followed by cosine annealing. See Appendix for details. After training, the model achieves a frame-level F1-score of around 91% on three test splits (Fig. 3). With a similar number of training samples, [gasp] is relatively difficult to perceive and detect, whereas [snore] and [yawn] are easier to distinguish due to their distinctive acoustic patterns. Similarly, [breath] performs slightly worse than [crying] and [throatclearing]. Overall, the frame-level detection model performs relatively poorly on [gasp] and [yawn] due to limited training samples, but performs well on other NV types. Notably, even when trained solely on English speech, the model performs well on Chinese test data, demonstrating strong language-agnostic capabilities and potential for scalability.
Figure 3: Frame-level detection model test results and distribution of training NV segments.
Table 2: Evaluation results of NV speech generation on English (EN) and Chinese (ZH) test sets.